OpenStudy (anonymous):

The Waltons, a family of 12, are being given a $330 government tax deduction. They are given $30 for each female and $20 for each male. Build a system of equations that can be used to find the number of females and males in the Walton family.

OpenStudy (bahrom7893):

30f+20m = 330m f+m = 12

OpenStudy (bahrom7893):

sorry typo

OpenStudy (bahrom7893):

30f+20m = 330 f+m = 12

OpenStudy (anonymous):

Thank you :-)